What five numbers average to 84
Answer by checkley77(12844)

5*84=420 ANS. FOR THE TOTAL OF THE 5 NUMBERS.
EXAMPLE:
PICK ANY FIVE NUMBERS THAT TOTAL 420.
76+80+84+88+92=420
420/5=84